Is the function f : R ➡R, defined by f(x) = 1-|x| is differentiable in its domain.

1
Expert's answer
2021-11-01T16:15:59-0400

Let us check whether the function "f : \R \to\R" , defined by "f(x) = 1-|x|" is differentiable in its domain.

Since "\lim\limits_{x\to 0+}\frac{f(x)-f(0)}{x-0}
=\lim\limits_{x\to 0+}\frac{1-|x|-1}{x}
=\lim\limits_{x\to 0+}\frac{-|x|}{x}
=\lim\limits_{x\to 0+}\frac{-x}{x}=-1" and

"\lim\limits_{x\to 0-}\frac{f(x)-f(0)}{x-0}
=\lim\limits_{x\to 0-}\frac{1-|x|-1}{x}
=\lim\limits_{x\to 0-}\frac{-|x|}{x}
=\lim\limits_{x\to 0-}\frac{-(-x)}{x}
=\lim\limits_{x\to 0-}\frac{x}{x}=1\ne -1,"

we conclude that the function "f" is not differentiable at the point "x=0" , and hence it is not differentiable in its domain.